# Which course? Discourse! Teaching Discourse and Generation in the Era of LLMs

### v1

@@ 1 Introduction
Natural Language Processing (NLP) as a field today is one filled with intensely rapid developments and clashing view points: “Scaling is powerful” Kaplan et al. (2020); Hoffmann et al. (2022) vs. “Scaling is not all you need” Li et al. (2025a); Marcus (2025); Hooker (2025); “LLMs cannot possibly tell us anything about language” Chomsky (2023); Bolhuis et al. (2024) vs. “Actually, they can” Piantadosi (2024); Futrell and Mahowald (2025); “LLMs cannot model meaning because they are not grounded” Bender and Koller (2020) vs. “But grounding is not always needed” Piantadosi and Hill (2022); Pavlick (2023); “LLMs can be proxies for human annotators” Gilardi et al. (2023); Calderon et al. (2025) vs. “No, that’s risky” Baumann et al. (2025); Wang et al. (2025), to name a few. These debates no longer focus on NLP alone; they recontextualize how we think about linguistics and many other disciplines. Thus, in order for the next generation of AI practitioners to grasp the cutting edge, to critically engage in these debates, and to contribute new insights, we believe it is crucial to create curricula sitting at the intersection of different (sub)areas.
@@ 3 Design Principles
We believe in engaging undergraduate students in active research, not only through individual research experiences, but also in the classroom. The fast-paced nature of today’s NLP landscape makes it especially important to do so, and we believe that exposing students to recent work that is core to discourse processing, yet not necessarily labeled so, is important in helping them recognize the interplay between theory and practice. These topics are either incorporated into course material organically, or delivered as invited talks from speakers throughout the semester. This also help students to develop an appreciation for rigorous dataset and experimental design in research.
@@ 4.2 Content on Discourse
Entities introduced throughout the discourse plays a central role in coherence. We introduce the classic Centering Theory Grosz et al. (1995) which tracks the focus of attention, as well as the more flexible Centering in Optimality Theory Beaver (2004). Students work through examples for both accounts in a workshop style, as described in Section 3. These concepts can be directly used to analyze and assess the coherence of a document. We introduce approaches inspired by Centering Theory, such as Entity Grid Barzilay and Lapata (2008) and DiscoScore Zhao et al. (2023). Conversely, generation can inform discourse: e.g., summarization informs entity salience Lin and Zeldes (2025).
@@ 5 Assignments & Course Project
The second assignment focuses on students running three prompting experiments: (1) implementing and comparing BLEU and LLM-as-a-judge Kocmi and Federmann (2023) for machine translation evaluation; (2) classifying discourse relations (a subset of the DISRPT 2025 shared task datasets, Braud et al. 2025, which cover a wide variety of languages and genres and are represented in a unified format across frameworks) using a LLM-as-a-judge paradigm; and (3) prompting LLMs to generate RST trees. These experiments (especially 2 and 3) are open-ended, providing students with the opportunity to practice prompt engineering using the various principles covered in lectures, as well as the opportunity to experiment with a variety of model sizes. Students were given basic boilerplate code to call models and set up basic formatting, as well as the data necessary to complete the assignment. Tasks 1 and 2 integrate non-English data, with Task 1 allowing students to bring in their own background by providing their own. In addition, students were asked to conduct an analysis given various aspects with regard to performance correlation with model size and choice of models, proposing ways to improve the performance, and the errors the model tends to make etc.
@@ 7 Challenges and Future Directions
In addition, our homework assignments were designed to be open-ended. This worked well and was positively received by students. However, such open-ended homework assignments did entail manual grading, which is difficult to scale up to a larger class. Innovative ways that combine the spirit of exploration in assignments and automatic grading is needed for scaling this to a larger class size.

### v2

@@ 1 Introduction
Natural Language Processing (NLP) as a field today is one filled with intensely rapid developments and clashing view points: “Scaling is powerful” Kaplan et al. (2020); Hoffmann et al. (2022) vs. “Scaling is not all you need” Li et al. (2025a); Marcus (2025); Hooker (2025); “LLMs cannot possibly tell us anything about language” Chomsky et al. (2023); Bolhuis et al. (2024) vs. “Actually, they can” Piantadosi (2024); Futrell and Mahowald (2025); “LLMs cannot model meaning because they are not grounded” Bender and Koller (2020) vs. “But grounding is not always needed” Piantadosi and Hill (2022); Pavlick (2023); “LLMs can be proxies for human annotators” Gilardi et al. (2023); Calderon et al. (2025) vs. “No, that’s risky” Baumann et al. (2025); Wang et al. (2025), to name a few. These debates no longer focus on NLP alone; they recontextualize how we think about linguistics and many other disciplines. Thus, in order for the next generation of AI practitioners to grasp the cutting edge, to critically engage in these debates, and to contribute new insights, we believe it is crucial to create curricula sitting at the intersection of different (sub)areas.
@@ 3 Design Principles
We believe in engaging undergraduate students in active research, not only through individual research experiences, but also in the classroom. The fast-paced nature of today’s NLP landscape makes it especially important to do so, and we believe that exposing students to recent work that is core to discourse processing, yet not necessarily labeled so, is important in helping them recognize the interplay between theory and practice. These topics are either incorporated into course material organically, or delivered as invited talks from speakers throughout the semester. This also helps students to develop an appreciation for rigorous dataset and experimental design in research.
@@ 4.2 Content on Discourse
Entities introduced throughout the discourse play a central role in coherence. We introduce the classic Centering Theory Grosz et al. (1995) which tracks the focus of attention, as well as the more flexible Centering in Optimality Theory Beaver (2004). Students work through examples for both accounts in a workshop style, as described in Section 3. These concepts can be directly used to analyze and assess the coherence of a document. We introduce approaches inspired by Centering Theory, such as Entity Grid Barzilay and Lapata (2008) and DiscoScore Zhao et al. (2023). Conversely, generation can inform discourse: e.g., summarization informs entity salience Lin and Zeldes (2025).
@@ 5 Assignments & Course Project
The second assignment focuses on students running three prompting experiments: (1) implementing and comparing BLEU and LLM-as-a-judge Kocmi and Federmann (2023) for machine translation evaluation; (2) classifying discourse relations (a subset of the DISRPT 2025 shared task datasets, Braud et al. 2025, which cover a wide variety of languages and genres and are represented in a unified format across frameworks) using an LLM-as-a-judge paradigm; and (3) prompting LLMs to generate RST trees. These experiments (especially 2 and 3) are open-ended, providing students with the opportunity to practice prompt engineering using the various principles covered in lectures, as well as the opportunity to experiment with a variety of model sizes. Students were given basic boilerplate code to call models and set up basic formatting, and the data necessary to complete the assignment. Tasks 1 and 2 integrate non-English data, with Task 1 allowing students to bring in their own background by providing their own. Students were also asked to conduct an analysis given various aspects with regard to performance correlation with model size and choice of models, proposing ways to improve the performance, and the errors the model tends to make etc.
@@ 7 Challenges and Future Directions
In addition, our homework assignments were designed to be open-ended. This worked well and was positively received by students. However, such open-ended homework assignments did entail manual grading, which is difficult to scale up to a larger class. Innovative ways that combine the spirit of exploration in assignments and automatic grading are needed for scaling this to a larger class size.